MINUTES OF MEETING

peletonwhoopΠολεοδομικά Έργα

26 Νοε 2013 (πριν από 3 χρόνια και 8 μήνες)

82 εμφανίσεις



Doc No: Mice Agenda
-
min
-
116

Issue: 1

Date: 8
th

December 2010

Page
1

of
4




MINUTES OF MEETING


Title/Subject
:

MICE Schedule Coordination, Update and Planning

Venue:

R66, G


〶†MAi

Chairman:

Willie Spensley

Time/Date:

13.30, 8
th

December, 2010

Secretary:

W.Spensley on behalf of Andy
Nichols

PARTICIPANTS


DISTRIBUTION

Matt Hills

W.Spensley

Geoff Barber

Mike Courthold

Craig Macwaters

David Pyke

Henry Nebrensky

Vishal Francis

Pierrick Hanlet


MICE Web Page

APOLOGIES

Andy Nichols

Ken.
Long



AGENDA REF:

Actions from last meeting

ATTACHMENTS:

None



1. Decay Solenoid & cryogenic systems


Matt Hills, Mike Courthold

Waiting for spare mini
-
compressor quote


Willie.
-

stands

John Govans to put signage on main air valves


stands,
John

has informed Willie that he
is placing signs & locking devices on valve handles

Willie


fit the serviced scroll pump. Stands


Done
, one remaining

Mike


write down RAMS for He
lium

buffer leak access for repair
.

Stands

& Blank
required

Automated cool
down software is loaded


Mike ask Pierrick to re
-
boot over the network.

Done

Cool down

underway


Saturday IOC shut down and then re
-
started Mike/Matt
-

New
controls working well

-

Vacuum Gauges are dirty and require clean also may add extra
vacuum gauges



Mike/Matt

Need to connect vac
uum

pump to absorber


Matt/Vicky
Done

Need to order KF50 fittings and seals
-

Matt

Add anything necessary to MOM walkabout checklist


Matt/Mike/Vicky

Done

LN2 ordered


Andy order BIP grade helium pack
???


Note on conven
tional magnets

Matt Littlefield to report back on D1 positional data to Matt H


some feed back received,
but not conclusive
-

stands

Problem w
ith remote control to Q3

Martin Hughes has re
-
set dip switch settings

Spare power supply for Q1
-
Q3 due next ye
ar
-

Martin Hughes

Software no longer ramps in steps


Brian Martlew/Pete Owens


Henry to chase



Doc No: Mice Agenda
-
min
-
116

Issue: 1

Date: 8
th

December 2010

Page
2

of
4



Testing of Magnets D1&D2, Q3 this Friday


Henry & Pierrick


Liquid Hydrogen system
-

Matt Hills

Still need to work out conductance of pipes when pump on roof



Mike


stands

Roger can help assemble transfer line support
-

stands

UPS battery/diesel generator back up discussion going on


batteries look cheaper and would
fit in the MCR


Matt verify then present to ISIS as favored option.

Unable to fit flame t
raps to vents due to pressure drop
-

Matt, formulate presentation to
ISIS/MICE Safety Group

and also inform David Matthewson.

Stands

A.S.Visited last week and everything ok no foreseen problems

Delivery next week Cryostat & Unit

Matt supplied drawing to
Willie

Awaiting quote for stainless steel unistrut


2. Target & Upstream beam line


Ken Long

Useful target stator re
-
design workshop held at DL


this topic will be added to the large
-
scale magnetic simulation studies.

T2 presently stripped for some
attention/refinement to anti
-
rotation guide.

Workshop held at Daresbury last week further discussions on size of wire? .51 square
section. Another workshop to be held at Daresbury 21
st

Dec 2010


3. Luminosity Monitor


Paul Soler

30mm of extra shielding
might be required, depending on MICE’s future running
conditions


Paul Soler will provide details nearer the time


4. Tracker Integration in Lab7


Geoff Barber

Craig/Mike


waiting for hoses, delivery should fit with US visit.

Total mess up with shared
service
hoses not ordered as thought


Craig chasing

Heater
-

needs to be powered (minus 5 outside) soon, Craig has ordered three phase
connectors, delivery next week

Gordon Stanford has arranged for the 3 phase to be done next week

New sniffer arrived, t
ry leak detecting again


Craig/Mike

Vacuum down on Cryo 3 needs to be replaced


5. Installation of Step IV


Tim Hayler
.

Magnetic simulation exercise has been established as a ‘project’


Tim/Mike organizing.

Vector Fields visited this week and has tak
en a module away.


6. Hall Infrastructure


Tim Hayler/David Pyke

KL trolley modifications for interim EMR are done.



Duplicate beam stopper power pack is done
require wiring diagram


Geoff


Need to know where we cut the RF floor/North mezzanine floor


must have drawings first



Tim


Quote to remove concrete steps in RF area due in one week


Willie to re
-
route water
services, Ian electrical services.





Doc No: Mice Agenda
-
min
-
116

Issue: 1

Date: 8
th

December 2010

Page
3

of
4



All large orders now placed for roof installations and NE rack mezzanine extension now
placed


good news.


Waiting for main door to be repainted and PPS gates to be touched up


David


Users must be careful on roof while handrail work in progress (and i
cy weather)


List of Infrastructure tasks for this FY

:

• New North Mezzanine extension for electronics racks: Framework due to be installed
w/c 6 Dec, mesh flooring will be the week before Christmas or the week after. Order gone
out
Work to commence nex
t week a meeting to be arranged with David, Kier, Willie
Thursday this week

• Security fence and handrail for the MICE roof: done, minor mods and incorporation of
small gate awaited


weather trouble at present
.
Fence complete but mods required to
meet wit
h Paul Wright


• Walkway from stairs
to the side roof:
-

in progress.

Installed, Landscaping to be done,
Gate required


• Walkway behind south shield wall at mezzanine level


changed to open mesh floor,
waiting for price,
stands


• Gas bottle rack outside

the MICE hall: Drawings issued in two weeks
.
Test hole
required, tender maybe before Christmas



• Vacuum enclosure on roof:
-

order placed
Kier on Monday


• Hydrogen stacks platform to support Hydrogen vents: Due to be installed late January


see above
re flame trap


7. Electrical Infrastructure


Ian Mullacrane


Substation 25 upgrade tender out today or tomorrow


yet more good news

Work has started on LH2 Controls installation
and making good progress

Local lighting under South mezzanine might need
reviewed


Alastair Sharp


8. MLCR, controls and networking, Online

M.Courthold/Craig Mac waters

Henry
-

talk to JSG about installation of KVM switch.

Move windows PC to staff network otherwise get rid of it


Henry


note visitors should be
equipped with a federal ID anyway, Rose can sort this out.

Some localised worry about rack sockets & cabling in the rack room


Craig/Henry
converge and relate to B
rian Martlew


in progress


9. PPS


Craig MacWaters

Almost commissioned, functional test, paperwork & admin remains


John/Craig.

Included above should be a simple ‘user guide’ for MICE.

Andy


discuss with MICE at large ownership of keys, etc


10. RF

Amplifier Refurbishment Progress & Daresbury
-

Willie Spensley/Nigel Rimmer



Doc No: Mice Agenda
-
min
-
116

Issue: 1

Date: 8
th

December 2010

Page
4

of
4



Fixed the Aluminium Plate

into the bottom of the Base Frame.

Machined a new Drive Shaft to replace the one that was bent.

Started to assemble the Anode and Grid Tuning Drive

Mechanisms.

The remaining

components, that were sent for cleaning
, will

hopefully be finished sometime
this week

11. Dates of next shutdown and job list

All
.

Long shutdown for moderator change


15
th
August 2010


February 2011


• Geoff


Beam stopper d
uplicate po
wer pack


Wiring diagram required Geoff to Chase


• Geoff


Air operation to Proton absorber


design done, Roger Geoff assemble


• Mike/Willie


Refurbish and install DS pumping rigs


one left to refurbish


• DL/Mike


DS quench detector rebuild
Next FY


• Vicky/Mike/DL


DS automated cool down project


cool down started


• Andy


initiate thermal balance study in R5.2
Stands


• Willie


check flow meter connection on chiller in catacombs
done


• Mike/Willie


HE buffer system leak repair
awaiting & blank required


Note that during the long shutdown (August 2010


February, 2011), the ISIS outer
synchrotron will be a CDM area. Adrian McFarland will be the Principal Contractor and all
risk assessm
ents and method statements should be submitted to him. He will also
arrange personnel induction.

This will not supersede the CDM arrangements for R5.2


12. Date of next meeting

Wednesday, 15
th December, 2010, 13:30 local time, G
-
06, R66, RAL.

13. A.O.B

Christmas ‘briefing’ for MIC
E
and friends around mid
-
Decembe
r